OpenStudy (anonymous):

Write y = 3/8 x - 2 in standard form where A, B, and C are integers whose greates common factor is 1. 3x - 8y = 2 3/8 x - y = 2 3x - 8y = 16 3x + 8y = 16 I think it's c

OpenStudy (anonymous):

= (3/8)x - 2 multiply through by 8: 8y = 3x - 2 rewrite: 3x - 8y = 2

OpenStudy (anonymous):

I picked that last time and got it wrong :/

OpenStudy (anonymous):

oh, i dunno :/ thats what i got too!

OpenStudy (texaschic101):

y = 3/8x - 2 (multiply by 8) 8y = 3x - 16 (subtract 3x) -3x + 8y = - 16 (multiply by -1 to make x positive) 3x - 8y = 16

OpenStudy (texaschic101):

you are correct...it is C

OpenStudy (anonymous):

Yay! I pretty much did everything except multiply by -1

OpenStudy (texaschic101):

the reason I multiplied by -1 is to make x positive....you do not want a negative x

OpenStudy (anonymous):

Thanls XD I didn't know that :)

OpenStudy (texaschic101):

no problem :)